A firm has a Cobb-Douglas production function given as q=AX1αX2β



a. Solve for the factor demand functions for labour (X1) and Capital (X2)



b. If the firms’ competitive output price is p find the wage rate



c. What is the share of the firm’s revenue paid to labour and capital?



d. If α=0.6, β=0.2 and A=1 find the LR labour and capital demand curve equations



4. Suppose a firm produces according to the production function Q = AL0.6K0.2, and faces wage rate ₵10, a rental cost of capital ₵5, and sells output at a price of ₵20.



a. Obtain and expression for the factor demand functions if A=1.



b. Compute the profit-maximizing factor demands for capital and labour if A=1.



5. In the short run, a competitive firm has a production function,



Q = f(L) = 2.6667L0.75. The output price is $4 per unit and the wage is $5 per hour. Find the short-



run labor demand curve of the firm.




1
Expert's answer
2021-11-03T10:39:30-0400

(a)

Factor demand functions for labor and capital

conditioned factor demand of capital:

"K=K(r,w,Q)"

where r=rate, w=wage and Q=Quantity.

To get the factor demand function of capital, we keep "w" and "Q" constant.

This results to:

"X_2=X_2(r)"

The conditional factor demand of labor:

"L=L(r,W,Q)"

where r=rate, W= wage and Q=Quantity

when we have "W" and "Q" as constants, the resulting labor force demand function is:

"X_1=X_1(r)" .


(b)

Wage rate = Marginal Productivity of Labor.

"MPL(Al\alpha X_2\beta)"

the wage rate will be: "A\alpha X_2\beta" .

(c)

Firm revenue paid to labor:

"=MPL\times Q"

but "MPL=A^\alpha X_2^\beta"

"\therefore MPL\times Q=(A\alpha X_2\beta)=Q(A\alpha X_2\beta)"

Firm revenue paid to capital :

"=MPK\times Q"

"=AX_1\alpha \beta\times Q"

"=Q(AX_1\alpha \beta )" .


(d)

"Q=AX_1\alpha X_2\beta"

The demand equation for labor "=(A\alpha X_2\beta)"

Substituting the values of "\alpha" ,"\beta" and "A" :

"Q=0.6K0.2=1.2K"


The demand equation for capital"=(AX_1 \alpha \beta)"

substituting the values of "\alpha" ,"\beta" and "A" :

"Q=X_1\times 0.6\times 0.2= 1.2X_1"